OpenStudy (anonymous):

A puppy pen is 1 foot longer than twice its width. John wants to increase the length and width by 5 feet each to enlarge the area by 90 square feet. What will be the area of the new pen?

OpenStudy (ranga):

Assume width = W Length = 2W + 1 (because “the length is 1 foot longer than twice its width”) Area = Length x Width = W(2W + 1) New Width = old Width + 5 = (W + 5) (because John increases the Width by 5 feet) New Length = old Length + 5 = (2W + 1) + 5 = (2W + 6) (because John increases the Length by 5 feet) New Area = New Length x New Width = (2W + 6)(W + 5) Increase in Area = New Area – old Area = (2W + 6)(W + 5) - W(2W + 1) = 90 (because the area increased by 90) Expand the expression and solve for W: (2W + 6)(W + 5) - W(2W + 1) = 90 Once you solve for W, plug the value into the equation for New Area, which is, (2W + 6)(W + 5) and you will get the answer to the question.